The ‘ze’ fonts (virtual T1 fonts) version 2.1

Robert Fuster

Departament de Matemàtica Aplicada Universitat Politècnica de València 46071 València (Spain)

                             e-mail: rfuster mat.upv.es
                         URL: HTTP://www.upv.es/~rfuster

June 24, 1997 September 22, 1998 July 17, 2000

Contents

1 Description 1

2 Installation 2

3 Warnings 2

1 Description

The ‘zd’ fonts by Constantin Kahn (kahn@math.uni-hannover.de) are virtual T1 encoded Computer Modern fonts based on (OT1) Computer Modern, Times, and Helvetica fonts, intended for simulate ‘dc’ fonts.^{1}

Because ‘dc’ fonts are now obsolete, I’ve adapted the Kahn’s package to ‘ec’ fonts. The resulting virtual fonts are named according to the ec fonts names, changing ‘ec’ by ‘ze’ (zerm1000.vf simulates ecrm1000, and so on)

Another virtual T1 encoded Computer Modern fonts are the almost european (‘am’) fonts, by Lars Engebretsen (enge@nada.kth.se). The main difference between ‘am’ and ‘ze’ fonts is that almost european fonts only use Computer Modern fonts to simulate the European Modern and, then, more characters are lost.

1Waine Sullivan’s ‘dm’ fonts are another approach to the substitution of ‘dc’ fonts by virtual ones.

2 Installation

The installation is very easy: you only need to put every file in its correct directory.

• Copy all the files from the vf directory where your dvi driver can find virtual files.^{2}

• Copy all the files from the tfm directory where TEX look for font metrics.

• Copy all the files from the texinput directory where L^{A}TEX look for input files.

• If you have the dvips driver, copy the files from the dvips directory where dvips looks for font mapping.

• If you have an (very) old version of zefonts, remove this files:

– cmsy10o.vf cmmi10o.vf cmvtt10o.vf cmbsy10o.vf cmmib10o.vf and cmbxi10f.vf (from your virtual fonts path),

– ze.map and config.ze (from the directory where dvips search for font mapping).

3 Warnings

• This is an unsupported software, and it comes without any kind of warranty.

• You can distribute it freely.

• My work for construct the ‘ze’ fonts was a simple adaptation from the ‘zd’ fonts by C. Khan.

2Your dvi driver must support virtual fonts, of course.